# Refined half-integer condition on RG flows

### v1

@@ 1 Introduction and summary
First, in a diagonal RCFTs, there is a one-to-one correspondence between primary operators and (zero-form) symmetry operators called Verlinde lines [18]. [^6] While there is a general proof of the correspondence, we find it more educational to look at a simple example. Consider the Yang-Lee model or $M(5,2)$ in the notation of the yellow book [19]. It is a diagonal RCFT with two primary operators with conformal dimensions $0,-\frac{1}{5}$. We place the theory on a two-torus with modular parameter $\tau$. Its modular $S$-matrix is given by (in the basis) $S=-\sqrt{\frac{2}{5-\sqrt{5}}}\begin{pmatrix}1&-\zeta^{-1}\\ -\zeta^{-1}&-1\end{pmatrix}\quad(\zeta:=\frac{1+\sqrt{5}}{2}).$ Since a symmetry operator is topological, it commutes with the energy-momentum tensor, and hence commutes with all Virasoro generators. Therefore, symmetry actions cannot change conformal dimensions. We learn (internal) symmetries act diagonally on the primaries. Let us denote a putative symmetry operator by $c$, and its action on the primaries as $a,b$, respectively. This should make it clear that the maximum number of (independent) symmetry transformations is two. In terms of partition function, if we insert the symmetry operator parallel to a ‘time’ slice, by shifting and acting it on the Hilbert space $\mathcal{H}$, we get the twisted partition function $Z^{c}(\tau,\bar{\tau}):=\text{tr}_{\mathcal{H}}\left(\hat{c}q^{L_{0}-c/24}\bar{q}^{\bar{L}_{0}-c/24}\right)=a|\chi_{0}(\tau)|^{2}+b|\chi_{-\frac{1}{5}}(\tau)|^{2},$ where $c=-\frac{22}{5}$. The RHS can be written as $Z^{c}(\tau,\bar{\tau})=\sum_{i,j}(M_{c})_{i,j}\chi_{i}(\tau)\bar{\chi}_{j}(\bar{\tau})$ with the mass matrix $M_{c}=\begin{pmatrix}a&0\\ 0&b\end{pmatrix}$. The actions $a,b$ cannot be arbitrary. It should obey the Cardy condition [20, 21]; since we can view the other direction as ‘time,’ the twisted partition function should have a physical interpretation in terms of (defect) Hilbert space. In equations, we perform the modular $S$-transformation. The symmetry operator $c$ is now inserted along the ‘time’ direction, and it intersects a ‘time’ slice. Thus, the original Hilbert space is modified to the defect Hilbert space $\mathcal{H}_{c}$. For the twisted partition function $Z_{c}(\tau,\bar{\tau}):=Z^{c}(-1/\tau,-1/\bar{\tau})=\text{tr}_{\mathcal{H}_{c}}\left(q^{L_{0}-c/24}\bar{q}^{\bar{L}_{0}-c/24}\right)=\sum_{i,j}(SM_{c}S^{\dagger})_{i,j}\chi_{i}(\tau)\bar{\chi}_{j}(\bar{\tau})$ (1.2) to have physical interpretation, all the coefficients have to be natural numbers $\forall i,j,\quad(SM_{c}S^{\dagger})_{i,j}\stackrel{{\scriptstyle!}}{{\in}}\mathbb{N}.$ (1.3) With the modular $S$-matrix and our mass matrix, this condition reduces to components of $\begin{pmatrix}A&B\\ B&C\end{pmatrix}$ with $A:=\frac{(5+\sqrt{5})a+(5-\sqrt{5})b}{10},B:=\frac{-a+b}{\sqrt{5}},C:=\frac{(5-\sqrt{5})a+(5+\sqrt{5})b}{10}$ be natural numbers. By adding $A$ and $C$, we get $a+b,\frac{-a+b}{\sqrt{5}}\in\mathbb{N}.$ Clearly, the identity operator $a=1=b$ satisfies the Cardy condition. Since the maximum number of symmetry operators is two, and we have already found one, we only have to fine another. For the other solution to the Cardy condition to make the second ratio a natural number, the difference of $a$ and $b$ must be a multiple of $\sqrt{5}$. Since we know $\zeta=\frac{1+\sqrt{5}}{2},-\zeta^{-1}=\frac{1-\sqrt{5}}{2}$, these numbers could work as $a$ and $b$. For the difference a+b-a+b to be positive, we have to take $a=-\zeta^{-1},\quad b=\zeta.$ Then, $a+b=1,\frac{-a+b}{\sqrt{5}}=1$, and the Cardy condition is satisfied. Since we have found two solutions, we can stop here. A general proof shows that a symemtry operator $c_{i}$ corresponding to a primary $\phi_{i}$ (with $\phi_{1}$ the identity operator) acts on a primary operator $\phi_{j}$ by $c_{i}\phi_{j}=\frac{S_{ij}}{S_{1j}}\phi_{j},$ (1.4) where $S_{ij}$ is the modular $S$-matrix. The braiding of Verlinde lines are characterized by the conformal dimensions of the corresponding primary operators. [^7] For example, a double braiding of $c_{i},c_{j}$ is given by $c_{c_{j},c_{i}}\circ c_{c_{i},c_{j}}\cong\bigoplus_{k=1}^{\text{rank}(\mathcal{C})}{N_{ij}}^{k}\frac{e^{2\pi ih_{k}}}{e^{2\pi i(h_{i}+h_{j})}}id_{k},$ where ${N_{ij}}^{k}\in\mathbb{N}$ is the fusion coefficient counting the number of (simple) object $c_{k}$ in the fusion $c_{i}\otimes c_{j}$, and $id_{k}$ is the identity morphism at $c_{k}$. Here, $\text{rank}(\mathcal{C})$ is the rank of $\mathcal{C}$ counting the number of simple objects in $\mathcal{C}$. Therefore, if a symmetry operator $c$ survives a relevant deformation, and if the RG flow ends up in another diagonal [^8] This is an assumption. However, many non-diagonal RCFTs can be described in terms of diagonal ones. For example, the three-state Potts model is non-diagonal as a representation of the Virasoro algebra, but it is diagonal as a representation of the extended $W_{3}$ algebra. In addition, many non-diagonal RCFTs are obtained by gauging anomaly-free (discrete) symmetries in diagonal RCFTs. If one is interested in RG flows of such non-diagonal RCFTs, one can study the flow from those in diagonal cases when the RG flows are triggered by operators invariant under the gauged symmetries employing the commutativity of relevant deformation and discrete gauging [10]. RCFT, it admits two conformal dimensions, one in UV $h_{c}^{\text{UV}}$ and the other in IR $h_{F(c)}^{\text{IR}}$, where $F:\mathcal{S}_{\text{UV}}\to\mathcal{S}_{\text{IR}}$ is the monoidal functor giving the RG flow. It was shown in [14, 15] that in RG flows with RG defects, their conformal dimensions should add up to half-integers (1.1). (Thus, the monoidal functor is not braided.) The RG defects [16] enter the proof as follows. As in our case, suppose a relevant deformation of a UV RCFT triggers an RG flow to an IR RCFT. By assumption, if we turn on the relevant coupling(s) in the whole two-dimensional space(time), the UV RCFT turns into the IR RCFT in the whole space(time). Now, let us turn on the relevant coupling(s) in only half of the space(time). Then, the other half remains the UV RCFT, while the deformed half flows to the IR RCFT. The system develops a defect between two theories. [^9] This picture makes it clear that the interface can be viewed as an object inserted where the relevant couplings abruptly jump from zero to non-zero values. Therefore, in the terminology of [17], we believe it is appropriate to call it an RG defect rather than RG domain wall or RG interface. The defect is called an RG defect (or RG domain wall/interface). Gaiotto constructed the defect for RG flows between diagonal cosets, however, he claimed the construction works more generally. In fact, the intuition does not seem to have any obstruction for such RG flows to admit RG defects. If an RG flow between two RCFTs admits an RG defect, we can employ the folding trick to place two RCFTs on the same half of the space(time). (An RG defect becomes an RG boundary condition.) Then, we can consider a product symmetry operator $cF(c)$. It has (chiral) disorder (a.k.a. twist) operator with conformal dimension $(h_{c}^{\text{UV}}+h_{F(c)}^{\text{IR}},0),(0,h_{c}^{\text{UV}}+h_{F(c)}^{\text{IR}})$ at the end as the product symmetry operator obeys the trivial fusion $cF(c)\otimes 1\cong cF(c)$. Gaiotto argued [16] that the disorder operators define extended vertex operator (super)algebra. Therefore, the sum has to be a half-integer (1.1). [^10] More precisely, the RG flow has to be simple. An RG flow is called simple [13] if it cannot be divided into more than one RG flows. On the other hand, a sequence of more than one RG flows is called non-simple. An example of non-simple RG flow is tetracritical Ising model $M(6,5)\to$ tricritical Ising model $M(5,4)\to$ critical Ising model $M(4,3)$. In order to get a non-simple RG flow, one has to deform with relevant operators more than once because one relevant deformation brings a UV RCFT to an IR RCFT. In other words, non-simple RG flows in general do not support nontrivial surviving symmetries throughout the sequence of RG flows. Concretely, if the first RG flow preserves $\mathcal{S}_{1}$ and the second $\mathcal{S}_{2}$, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ do not have to have common symmetry objects except the identity object. When one tries to solve an RG flow, one performs relevant deformation only once. Thus, the flows are believed to be simple.
@@ 2 Examples
for a Cardy state $|a\rangle$ with variational parameter $\tau_{a}$. Here, $j$ is the index corresponding to the relevant operator $\phi_{\frac{29}{31}}$. Tha ratio of the modular $S$-matrices takes the largest value ($\approx 2.959$) for Cardy states with conformal dimensions $0,\frac{47}{5}$, and the smallest value ($\approx-0.990$) for those with $\frac{99}{31},\frac{92}{155}$. Therefore, for positive Lagrangian couplings, the first two Cardy states give (approximate) vacuum, and for negative Lagrangian couplings, the second two Cardy states give twofold degenerate vacua. Since the first Cardy states include the identity, this signals the massless flow. To summarize, we arrive the conjecture
@@
where the two Fibonacci objects both have conformal dimensions $h_{\text{Fib}}=\frac{23}{5}$. We deform the theory with the least relevant operator with conformal dimension $\frac{9}{11}$. It does not turn on other relevant operators because it obyes the OPE
@@
where we denote $\mathbb{Z}/3\mathbb{Z}$ objects in UV by $x,x^{2}$, and their images in IR by $X,X^{2}$. We have $h_{X}=\frac{2}{3}=h_{X^{2}}$. This concludes our analysis for the massless case. For the massive case, we know the IR TQFT is described by a $\text{Vec}_{\mathbb{Z}/3\mathbb{Z}}^{1}$-module category. Since the surviving symmetry is an anomaly-free $\mathbb{Z}/3\mathbb{Z}$ symmetry, both one- and three-dimensional modules (i.e., representations) exist:

### v2

@@ 1 Introduction and summary
First, in a diagonal RCFTs, there is a one-to-one correspondence between primary operators and (zero-form) symmetry operators called Verlinde lines [18]. [^6] While there is a general proof of the correspondence, we find it more educational to look at a simple example. Consider the Yang-Lee model or $M(5,2)$ in the notation of the yellow book [19]. It is a diagonal RCFT with two primary operators with conformal dimensions $0,-\frac{1}{5}$. We place the theory on a two-torus with modular parameter $\tau$. Its modular $S$-matrix is given by (in the basis) $S=-\sqrt{\frac{2}{5-\sqrt{5}}}\begin{pmatrix}1&-\zeta^{-1}\\ -\zeta^{-1}&-1\end{pmatrix}\quad(\zeta:=\frac{1+\sqrt{5}}{2}).$ Since a symmetry operator is topological, it commutes with the energy-momentum tensor, and hence commutes with all Virasoro generators. Therefore, symmetry actions cannot change conformal dimensions. We learn (internal) symmetries act diagonally on the primaries. Let us denote a putative symmetry operator by $c$, and its action on the primaries as $a,b$, respectively. This should make it clear that the maximum number of (independent) symmetry transformations is two. In terms of partition function, if we insert the symmetry operator parallel to a ‘time’ slice, by shifting and acting it on the Hilbert space $\mathcal{H}$, we get the twisted partition function $Z^{c}(\tau,\bar{\tau}):=\text{tr}_{\mathcal{H}}\left(\hat{c}q^{L_{0}-c/24}\bar{q}^{\bar{L}_{0}-c/24}\right)=a|\chi_{0}(\tau)|^{2}+b|\chi_{-\frac{1}{5}}(\tau)|^{2},$ where $c=-\frac{22}{5}$. The RHS can be written as $Z^{c}(\tau,\bar{\tau})=\sum_{i,j}(M_{c})_{i,j}\chi_{i}(\tau)\bar{\chi}_{j}(\bar{\tau})$ with the mass matrix $M_{c}=\begin{pmatrix}a&0\\ 0&b\end{pmatrix}$. The actions $a,b$ cannot be arbitrary. It should obey the Cardy condition [20, 21]; since we can view the other direction as ‘time,’ the twisted partition function should have a physical interpretation in terms of (defect) Hilbert space. In equations, we perform the modular $S$-transformation. The symmetry operator $c$ is now inserted along the ‘time’ direction, and it intersects a ‘time’ slice. Thus, the original Hilbert space is modified to the defect Hilbert space $\mathcal{H}_{c}$. For the twisted partition function $Z_{c}(\tau,\bar{\tau}):=Z^{c}(-1/\tau,-1/\bar{\tau})=\text{tr}_{\mathcal{H}_{c}}\left(q^{L_{0}-c/24}\bar{q}^{\bar{L}_{0}-c/24}\right)=\sum_{i,j}(SM_{c}S^{\dagger})_{i,j}\chi_{i}(\tau)\bar{\chi}_{j}(\bar{\tau})$ (1.2) to have physical interpretation, all the coefficients have to be natural numbers $\forall i,j,\quad(SM_{c}S^{\dagger})_{i,j}\stackrel{{\scriptstyle!}}{{\in}}\mathbb{N}.$ (1.3) With the modular $S$-matrix and our mass matrix, this condition reduces to components of $\begin{pmatrix}A&B\\ B&C\end{pmatrix}$ with $A:=\frac{(5+\sqrt{5})a+(5-\sqrt{5})b}{10},B:=\frac{-a+b}{\sqrt{5}},C:=\frac{(5-\sqrt{5})a+(5+\sqrt{5})b}{10}$ be natural numbers. By adding $A$ and $C$, we get $a+b,\frac{-a+b}{\sqrt{5}}\in\mathbb{N}.$ Clearly, the identity operator $a=1=b$ satisfies the Cardy condition. Since the maximum number of symmetry operators is two, and we have already found one, we only have to fine another. For the other solution to the Cardy condition to make the second ratio a natural number, the difference of $a$ and $b$ must be a multiple of $\sqrt{5}$. Since we know $\zeta=\frac{1+\sqrt{5}}{2},-\zeta^{-1}=\frac{1-\sqrt{5}}{2}$, these numbers could work as $a$ and $b$. For the difference a+b-a+b to be positive, we have to take $a=-\zeta^{-1},\quad b=\zeta.$ Then, $a+b=1,\frac{-a+b}{\sqrt{5}}=1$, and the Cardy condition is satisfied. Since we have found two solutions, we can stop here. A general proof shows that a symmetry operator $c_{i}$ corresponding to a primary $\phi_{i}$ (with $\phi_{1}$ the identity operator) acts on a primary operator $\phi_{j}$ by $c_{i}\phi_{j}=\frac{S_{ij}}{S_{1j}}\phi_{j},$ (1.4) where $S_{ij}$ is the modular $S$-matrix. The braiding of Verlinde lines are characterized by the conformal dimensions of the corresponding primary operators. [^7] For example, a double braiding of $c_{i},c_{j}$ is given by $c_{c_{j},c_{i}}\circ c_{c_{i},c_{j}}\cong\bigoplus_{k=1}^{\text{rank}(\mathcal{C})}{N_{ij}}^{k}\frac{e^{2\pi ih_{k}}}{e^{2\pi i(h_{i}+h_{j})}}id_{k},$ where ${N_{ij}}^{k}\in\mathbb{N}$ is the fusion coefficient counting the number of (simple) object $c_{k}$ in the fusion $c_{i}\otimes c_{j}$, and $id_{k}$ is the identity morphism at $c_{k}$. Here, $\text{rank}(\mathcal{C})$ is the rank of $\mathcal{C}$ counting the number of simple objects in $\mathcal{C}$. Therefore, if a symmetry operator $c$ survives a relevant deformation, and if the RG flow ends up in another diagonal [^8] This is an assumption. However, many non-diagonal RCFTs can be described in terms of diagonal ones. For example, the three-state Potts model is non-diagonal as a representation of the Virasoro algebra, but it is diagonal as a representation of the extended $W_{3}$ algebra. In addition, many non-diagonal RCFTs are obtained by gauging anomaly-free (discrete) symmetries in diagonal RCFTs. If one is interested in RG flows of such non-diagonal RCFTs, one can study the flow from those in diagonal cases when the RG flows are triggered by operators invariant under the gauged symmetries employing the commutativity of relevant deformation and discrete gauging [10]. RCFT, it admits two conformal dimensions, one in UV $h_{c}^{\text{UV}}$ and the other in IR $h_{F(c)}^{\text{IR}}$, where $F:\mathcal{S}_{\text{UV}}\to\mathcal{S}_{\text{IR}}$ is the monoidal functor giving the RG flow. It was shown in [14, 15] that in RG flows with RG defects, their conformal dimensions should add up to half-integers (1.1). (Thus, the monoidal functor is not braided.) The RG defects [16] enter the proof as follows. As in our case, suppose a relevant deformation of a UV RCFT triggers an RG flow to an IR RCFT. By assumption, if we turn on the relevant coupling(s) in the whole two-dimensional space(time), the UV RCFT turns into the IR RCFT in the whole space(time). Now, let us turn on the relevant coupling(s) in only half of the space(time). Then, the other half remains the UV RCFT, while the deformed half flows to the IR RCFT. The system develops a defect between two theories. [^9] This picture makes it clear that the interface can be viewed as an object inserted where the relevant couplings abruptly jump from zero to non-zero values. Therefore, in the terminology of [17], we believe it is appropriate to call it an RG defect rather than RG domain wall or RG interface. The defect is called an RG defect (or RG domain wall/interface). Gaiotto constructed the defect for RG flows between diagonal cosets, however, he claimed the construction works more generally. In fact, the intuition does not seem to have any obstruction for such RG flows to admit RG defects. If an RG flow between two RCFTs admits an RG defect, we can employ the folding trick to place two RCFTs on the same half of the space(time). (An RG defect becomes an RG boundary condition.) Then, we can consider a product symmetry operator $cF(c)$. It has (chiral) disorder (a.k.a. twist) operator with conformal dimension $(h_{c}^{\text{UV}}+h_{F(c)}^{\text{IR}},0),(0,h_{c}^{\text{UV}}+h_{F(c)}^{\text{IR}})$ at the end as the product symmetry operator obeys the trivial fusion $cF(c)\otimes 1\cong cF(c)$. Gaiotto argued [16] that the disorder operators define extended vertex operator (super)algebra. Therefore, the sum has to be a half-integer (1.1). [^10] More precisely, the RG flow has to be simple. An RG flow is called simple [13] if it cannot be divided into more than one RG flows. On the other hand, a sequence of more than one RG flows is called non-simple. An example of non-simple RG flow is tetracritical Ising model $M(6,5)\to$ tricritical Ising model $M(5,4)\to$ critical Ising model $M(4,3)$. In order to get a non-simple RG flow, one has to deform with relevant operators more than once because one relevant deformation brings a UV RCFT to an IR RCFT. In other words, non-simple RG flows in general do not support nontrivial surviving symmetries throughout the sequence of RG flows. Concretely, if the first RG flow preserves $\mathcal{S}_{1}$ and the second $\mathcal{S}_{2}$, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ do not have to have common symmetry objects except the identity object. When one tries to solve an RG flow, one performs relevant deformation only once. Thus, the flows are believed to be simple.
@@ 2 Examples
for a Cardy state $|a\rangle$ with variational parameter $\tau_{a}$. Here, $j$ is the index corresponding to the relevant operator $\phi_{\frac{29}{31}}$. The ratio of the modular $S$-matrices takes the largest value ($\approx 2.959$) for Cardy states with conformal dimensions $0,\frac{47}{5}$, and the smallest value ($\approx-0.990$) for those with $\frac{99}{31},\frac{92}{155}$. Therefore, for positive Lagrangian couplings, the first two Cardy states give (approximate) vacuum, and for negative Lagrangian couplings, the second two Cardy states give twofold degenerate vacua. Since the first Cardy states include the identity, this signals the massless flow. To summarize, we arrive the conjecture
@@
where the two Fibonacci objects both have conformal dimensions $h_{\text{Fib}}=\frac{23}{5}$. We deform the theory with the least relevant operator with conformal dimension $\frac{9}{11}$. It does not turn on other relevant operators because it obeys the OPE
@@
where we denote $\mathbb{Z}/3\mathbb{Z}$ objects in UV by $x,x^{2}$, and their images in IR by $X,X^{2}$. We have $h_{X}=\frac{2}{3}=h_{X^{2}}$. This concludes our analysis for the massless case. For the massive case, we know the IR TQFT is described by a $\text{Vec}_{\mathbb{Z}/3\mathbb{Z}}^{1}$-module category. [^25] The classification of indecomposable $\text{Vec}_{\mathbb{Z}/3\mathbb{Z}}^{1}$-module categories have been performed in [39, 52] employing the method developed in [38, 39, 40, 53, 52]. The set of indecomposable $\mathcal{C}$-module categories stands in bijection with another set of Lagrangian algebras in $\mathcal{Z}(\mathcal{C})$ for a fusion category $\mathcal{C}$ [30]. In [52], it was found that $\mathcal{Z}(\text{Vec}_{\mathbb{Z}/3\mathbb{Z}}^{1})$ has two Lagrangian algebras, thus two indecomposable $\text{Vec}_{\mathbb{Z}/3\mathbb{Z}}^{1}$-module categories. The indecomposable module categories were not identified in the paper, but it is not difficult to identify them. First, any fusion category has itself as an indecomposable module category called regular module category. Second, since $\mathbb{Z}/3\mathbb{Z}$ is anomaly-free, we know there exists rank one module category. Therefore, the two indecomposable $\text{Vec}_{\mathbb{Z}/3\mathbb{Z}}^{1}$-module categories have ranks three and one. Since the surviving symmetry is an anomaly-free $\mathbb{Z}/3\mathbb{Z}$ symmetry, both one- and three-dimensional modules (i.e., representations) exist: